OpenStudy (anonymous):

sin theta if csc theta =-2

OpenStudy (xapproachesinfinity):

Recall that sinx=1/cscx

OpenStudy (xapproachesinfinity):

so if \(\rm csc\theta=-2 \longleftrightarrow sin\theta=-1/2\)

OpenStudy (xapproachesinfinity):

\(\rm \color{navy}{\bigstar Welcome~to~OpenStudy!\bigstar}\)